What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 39 ?
Answer by stanbon(75887)

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 39 ?
---
a(6) = 4 + 5d = 39
5d = 35
d = 7
---------
a(32) = 4 + (32-1)7
-----
a(32) = 4 + 31*7
a(32) = 221
===========
Cheers,
Stan H.
===========